Find the other angle and sides of the triangle:

  1. Right triangle ABC with angle b=38° of side a=14 cm.
  2. Obtuse triangle DEF with sides d=12, e=16.3 cm, f=13.5 cm
1
Expert's answer
2021-01-08T14:35:31-0500

QUESTION 1


Step 1: We know from the the law of sines

(a / sin A) = (b / sin B) = (c / sin C)

Therefore:

(b / sin B) = (a / sin A)

(b / sin 380)= (14 / sin 520)

b = (sin 380 x 14) / sin 520

b = 10.94 cm

Step 2: From pythogras theorem we know:

A+ B= C2

142 + 10.942 = C2

315.68 = C2

C = 17.77 cm

Step 3: The interor angles of a triangle add up to 1800

900 + 380 + A = 1800

A = 1800 – 1280

A = 520


A = 14, B= 10.94, and C = 17.77

Angle <A = 52<B = 90 <C = 38


QUESTION 2


Step 1: Given the lengths of all three sides of any triangle, each angle can be calculated using the following equations:


e = 16.3 cm

f = 13.5 cm 

d = 12 cm

<D = arcCos [(e+ f– d2) /(2ef)] ... (eq i)

<E = arcCos [(d+ f– e2) /(2df)] ...(eq ii)

<F = arcCos [(d+ e– d2) /(2de)] ...(eq iii)


Step 2: Substitute the values to the equations i, ii, and iii


<D = arcCos [(16.3+ 13.5– 122) / (2 x 16.3 x 13.5)]

<D = 46.320

<E = arcCos [(13.5+ 12– 16.32) /(2 x 13.5 x 12)]

< E = 79.230

<F = arcCos [(16.3+ 12– 13.5) /(2 x 16.3 x 13.5)]

< F = 54.450


Therefore:

Angles <D = 46.32 <E= 79.230, and <F = 54.450

Side lengths e = 16.3 cm , f = 13.5 cm , and d = 12 cm